What is the moment of inertia of an object about an axis if there are three point masses (m1 = 1.5 kg, m2 = 1.75 kg, m3 = 2 kg) located at r1 = 0.10 m, r2 = 0.20 m, and r3 = 0.35 m distance from the axis of rotation?

Options:

A) 2.24 kgm2

B) 29.46 kgm2

C) 1.20 kgm2

D) 0.33 kgm2
Answer #1

moment of Inertia is given by = I = m * r2

where m = mass of the object

and r = distance of object from the axis of rotation

in this question

I = m1 * r12 + m2 * r22 + m3 * r32  

I = ( 1.5 * 0.12 ) + ( 1.75 * 0.22 ) + ( 2 * 0.352 )

I = 0.33 kg - m2

answer option D